please help, it is substitution: Last year you mowed grass and shoveled snow for 12 households.You earned $225 for mowing lawn last year and you earned $200 for shoveling snow. You earned a total of $2600 last year. How many households did you mow for? How many households did you shovel?

Answer by ankor@dixie-net.com(22740)

Last year you mowed grass and shoveled snow for 12 households.
You earned $225 for mowing lawn last year and you earned $200 for shoveling snow.
You earned a total of $2600 last year.
How many households did you mow for? How many households did you shovel?
:
let m = no. of households mowed for
let s - no. of households shoveled for
:
Write an equation for each statement
"Last year you mowed grass and shoveled snow for 12 households."
m + s = 12
s = (12-m); use this form for substitution
:
You earned $225 for mowing lawn last year and you earned $200 for shoveling snow. You earned a total of $2600 last year.
225m + 200s = 2600
Substitute (12-m) for s
225m + 200(12-m) = 2600
distribute
225m + 2400 - 200m = 2600
225m - 200m = 2600 - 2400
25m = 200
m = 200/25
m = 8 households mowed
and
s = 12 - 8
s = 4 households shoveled
:
:
Check this in the 2nd equation
225(8) + 200(4) =
1800 + 800 = 2600